OpenStudy (anonymous):

A pair of jeans cost $48. If the sales tax is 6.5%, what is the total cost of the item?

jimthompson5910 (jim_thompson5910):

what is 6.25% of $48 ?

OpenStudy (anonymous):

why 6.25? wouldnt it be 6.5?

jimthompson5910 (jim_thompson5910):

oh sorry misread

jimthompson5910 (jim_thompson5910):

yes 6.5% not 6.25%

jimthompson5910 (jim_thompson5910):

6.5% of 48 = 0.065*48 = ??

OpenStudy (anonymous):

then it would be 312.000

jimthompson5910 (jim_thompson5910):

3.12 actually

jimthompson5910 (jim_thompson5910):

you forgot to move the decimal to the left 2 spots

jimthompson5910 (jim_thompson5910):

this is the amount of tax you have to pay, so add this to the cost to get 48+3.12 = 51.12 so that's your total cost

OpenStudy (anonymous):

how did i forget to move left 2 spots?

OpenStudy (anonymous):

just wondering

jimthompson5910 (jim_thompson5910):

one way to do this is to do 65 times 48 to get 3120 then you have to move the decimal 3 spots to the left (because it's 0.065 and not 65) to go from 3120 to 3.12

jimthompson5910 (jim_thompson5910):

not 100% sure why you forgot that step though, it's ok

OpenStudy (anonymous):

oh ok i see. i did it another way and i see why i messed up. Thanks!

jimthompson5910 (jim_thompson5910):

ok great, glad it's all clicking now yw